OpenStudy (anonymous):

Can someone please show me the steps for 14^0. Working on a project and needs the steps to do it.

OpenStudy (alexandervonhumboldt2):

14^0=1 any number in 0 exponent is 1

OpenStudy (anonymous):

It equals 1

OpenStudy (alexandervonhumboldt2):

here is the proof: a^0=a^1/a^1 a^1/a^1=a/a=1

OpenStudy (alexandervonhumboldt2):

=> a^0=1

OpenStudy (anonymous):

thankyou so kindly everyone :)

OpenStudy (alexandervonhumboldt2):

np